OpenStudy (anonymous):

Calculus and Parametric Equations: An object is moving along a curve in the xy-plane has position (x(t), y(t)) at time t with dx/dt = 2sin(t^3) and dy/dt = cos(t^2) for 0 <= t <= 4. At time t = 1, the object is at the position (3,4). Find the position of the object at time t = 2.

OpenStudy (anonymous):

not sure are you supposed to integrate \(\frac{dy}{dx}\)?

OpenStudy (anonymous):

that would be awesome, but I'm not sure how to do that while dy/dx is still in terms of t..

OpenStudy (anonymous):

\[\frac{dy}{dx}=\frac{\frac{dy}{dt}}{\frac{dx}{dt}}\]

OpenStudy (anonymous):

oh i see the problem it is still in terms of \(t\)

OpenStudy (anonymous):

yup :)

OpenStudy (anonymous):

I wish both dy/dt and dx/dt had simple solutions.. but in this case they're not, so I think there's another way to do it using parametrics..

OpenStudy (anonymous):

yeah, you are not going to find an anti derivative of that anyway hmm i wonder what you can do? i would bump this question up, maybe some smart person can answer it

OpenStudy (anonymous):

bumped.. thanks. I know you're one of the smartest/most helpful people around here since I've been using openstudy for some years now and you're one of the people who could always answer my questions. :)

OpenStudy (anonymous):

You going to close this? You still stuck?

OpenStudy (anonymous):

still stuck, yes

OpenStudy (anonymous):

I think you should leave the answer in terms of definite integrals.

OpenStudy (anonymous):

\[ \mathbf r(2) = \left\langle \int_{1}^{2}\cos(t^2)\;dt + 3, \int_{1}^{2}2\sin(t^3)\;dt + 4 \right\rangle \]

OpenStudy (anonymous):

There isn't any particular theorem here that I know of that will make these integrals closed.

OpenStudy (anonymous):

Thanks, well I can carry out numeric integration using my TI-84.. and doing that seems to yield the correct answer (well only that you switched the places of the cos and sin terms [x is sin and y is cos in the prob]). The answer turns out to be (3.437, 3.557)

OpenStudy (anonymous):

Could you please explain to me the thought process of how you set up those integrals?

OpenStudy (anonymous):

\[ \mathbf r(b) - \mathbf r(a) = \int_a^b\mathbf r'(t)\;dt \]

OpenStudy (anonymous):

\(a=1\), \(b=2\)

OpenStudy (anonymous):

\[ \mathbf r(2) = \int_1^2\mathbf r'(t)\;dt+\mathbf r(1) \]

OpenStudy (anonymous):

Thanks, that confirms what I've thought. Thanks so much!